Hello. Can you please help me?
I need to make v the subject of the following formula (Einstein's theory of relativity). It's a fraction involving m followed by a little O as numerator, with a radical in the denominator.
m= (mo)/ (radical around the following:) 1-(v/c)^2
Thank you so much!

Answer by edjones(8007)

Solve for v.
m=(Mo)/sqrt( 1-(v^2/c^2))
m*sqrt( 1-(v^2/c^2))=Mo
sqrt( 1-(v^2/c^2))=Mo/m
1-(v^2/c^2)=Mo^2/m^2
v^2/-c^2=(mo^2/m^2)-1
v^2=-c^2((mo^2/m^2)-1)
v=sqrt(m^2-Mo^2)*|c/m|
.
Ed